Use the difference quotient (f(x+4)=f(x) on the following function. f(x) = 3x2 – 4
Answer #1

we have f(x) = ax²-4 Replace a by ath in f(x), we get NOW f(x+a)= 3(x+67-4 31&toxh+h)-4 = 3x²+62h + 36²-4 subtract fled from